OpenStudy (anonymous):

Two concentric spheres have radii of 5" and 6". Find the volume of the space between them.

OpenStudy (anonymous):

find the volume of the larger radius circle, and subtract it from the volume of the lower

OpenStudy (anonymous):

how do you do that?

OpenStudy (anonymous):

(4/3)(pi)(r^3)= volume

OpenStudy (anonymous):

still dont get it at all

OpenStudy (anonymous):

i suck at math lol

OpenStudy (whpalmer4):

\[V=(4/3){\pi}r^3\] \[V _{6"} = (4/3)\pi(6)^3=(4/3)\pi(216)=288\pi\] \[V _{5"} = (4/3)\pi(5)^3=(4/3)\pi(125)=(500/3)\pi\] \[V _{between} = V _{6"} - V _{5"} = \]

OpenStudy (whpalmer4):

Hint: \[288\pi = (3*288\pi)/3\]